OpenStudy (anonymous):

write the equation of the line that is parallel to the line whose equation is y = -4 and that passes through the point (1,5). a. y = -5 b. y = -4 c. y = 1 d. y = 5

OpenStudy (anonymous):

-4? b

OpenStudy (zale101):

y = -4 y=0x-4 Now, Using point slope formula, the slope is zero so m =0 x1,y1= (1,5) \(y-y_1=m(x-x_1)\) \(y-5=0(x-1)\)

OpenStudy (zale101):

with parallel lines, slopes have the same slope, but since the slope of y=-4 is zero, then the line that is parallel to it has the same slope which is zero.

OpenStudy (zale101):

Makes sense?

OpenStudy (anonymous):

yea but wat do i get for my answer???? its not -4??

OpenStudy (zale101):

Yeah, if you solve this \(y-5=0(x-1)\) you'll get the solution.

OpenStudy (zale101):

so, it is not -4

OpenStudy (zale101):

Can you solve this equation and solve for y?

OpenStudy (anonymous):

wats x?

OpenStudy (zale101):